OpenStudy (anonymous):

The length of a rectangle is 2 units and its width is square root of 2 unit. Is the area of the rectangle rational or irrational?

OpenStudy (campbell_st):

you will need to multiply length by width... \[A = 2 \times \sqrt{2}\] do you know the difference between rational and irrational...

OpenStudy (anonymous):

Irrational has decimals, e.g. 2.3453 Rational is a whole number, e.g. 2

OpenStudy (campbell_st):

so what would \[2 \sqrt{2} \] be... decimal or whole

OpenStudy (anonymous):

So the answer is irrational because \[2\sqrt{2} = 2.828487...\]